Circular No. 6375

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996N IN NGC 1398 S. D. Van Dyk, University of California at Berkeley; R. A. Sramek,

National Radio Astronomy Observatory (NRAO); K. W. Weiler and M. J. Montes, Naval Research Laboratory; and N. Panagia, Space Telescope Science Institute and European Space Agency, report: “Observations of SN 1996N made on Apr. 1.96 UT with the NRAO’s Very Large Array telescope show a probable detection at wavelength 3.6 cm (8.4 GHz) of a source with a flux density of 0.15 ± 0.03 mJy at α = 3^{h}38^{m}55^{s}.374, δ = −26^{o}20′03′′.96 (equinox J2000.0; ± 0′′.2 in each coordinate). This is coincident to within 1′′ with an optical position for the supernova of α = 3^{h}38^{m}55^{s}.31, δ = −26^{o}20′04′′.1, measured by Van Dyk from a B-band image made on Apr. 2.02 by P. Ugarte and N. Suntzeff at Cerro Tololo. Since SN 1996N was reported to be apparently two weeks past optical maximum on Mar. 23 (IAUC 6351), we are very likely detecting optically thin radio emission (see Weiler et al. 1986, Ap.J. 301, 790). Radio monitoring is underway at this and other wavelengths.”

GRO J1744–28 J. E. Morrison and S. Röser, Astronomisches-Rechen-Institut, Heidel-

berg, write: “Cole et al. (IAUC 6356) report on positions of possible optical counterparts for the radio source GRO J1744–28 based on HST-GSC reference positions. They find an offset of 0′′.9 between the radio coordinates of GRO J1744–28 and their star ‘a’ or ‘C’. They discuss the possibility of a GSC–PPM systematic offset, but find only −0′′.333 in α and −0′′.024 in δ, insufficient to account for the 0′′.9 radio-optical offset. We have completed GSC 1.2, a new astrometric reduction of GSC 1.1 onto the PPM system removing plate and magnitude dependent systematics. In a small field around the position of the radio source, we find 13 GSC stars between 17^{h}44^{m} and 17^{h}45^{m}, and −28^{o}50′ and −28^{o}40′, for which we have α(GSC1.2–GSC1.1) = −0′′.038 and δ(GSC1.2–GSC1.1) = +0′′.45. With these corrections, the offset between the radio position and the infrared position of ‘C’ reduces to 0′′.5, (radio–‘C’ = +0′′.18 in α, +0′′.45 in δ), which may possibly be caused by the offset between the PPM and the radio system. This strengthens the hypothesis that ‘C’ in the infrared and ‘a’ in the optical are the counterparts of GRO J1744–28.”

1996 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green